# NonlinearIncidence.bib

### v4

@@ Abstract
We expand the calculation of the upper bound on epidemic overshoot in SIR models to account for nonlinear incidence. We lay out the general procedure and restrictions to perform the calculation analytically. We demonstrate the procedure by working through several examples and also numerically study what happens to the upper bound on overshoot when nonlinear incidence manifests in the form of epidemic dynamics over a contact network. We find that both steeper incidence terms and larger contact heterogeneity can increase the range of communicable diseases at which the overshoot remains a relatively large public health hazard.
@@ Introduction
Compartmental models have been an invaluable tool for analyzing the dynamics of epidemics for the last century. In particular, the susceptible-infected-recovered (SIR) model has long been a workhorse compartmental model for describing transient epidemics due to its relative simplicity and has received a lot of attention in both the academic literature and the public health arena [anderson_infectious_1992, diekmann_mathematical_2013]. The mechanism of transmission of a communicable disease underlying this model and a variety of other contagion models is contact between healthy and infectious individuals, which results in conversion of the healthy individuals into an infected state. Choosing how to precisely define the transmission interaction between healthy and infectious individuals leads to a variety of models depending on the assumptions made. For instance, in a spatially-explicit model, one might represent the contact between individual members of the population by means of a network. Alternatively, if one is willing to assume all individuals within a compartment are identical, one can formulate a model using ordinary differential equations (ODEs) by assuming an incidence term for how the susceptible and infected compartments mix and generate new infected individuals.
@@
Overshoot is a concept from mathematics and control theory that quantifies the amount of excess from when a function exceeds its target value [doyle_feedback_2013, sethi_what_2019]. This concept has been applied to a variety of contexts ranging from ecological and environmental problems that pertain to overconsumption and sustainability [catton_overshoot_1982, wackernagel_tracking_2002, bradshaw_underestimating_2021, fanning_social_2022] to biotechnology that controls blood flow [rosengarten_overshoot_2002]. In the epidemiological context, the overshoot quantifies the number of individuals that become infected after the prevalence peak of infections occurs. In simple epidemics, as the peak of the epidemic coincides with the threshold at which transmission is sufficiently reduced so that the epidemic is no longer growing, the overshoot reflects the excess in cases beyond this minimal threshold of protection.
@@
While the overshoot within the SIR model has received some attention [handel_what_2006, ellison_implications_2020, rachel_analytical_2020, ketcheson_optimal_2021], a detailed understanding of its full mathematical behavior within compartmental models remains incomplete. An observation that the overshoot is largest for intermediate basic reproduction numbers was first numerically observed by Zarnitsyna et al. [zarnitsyna_intermediate_2018]. An explanation for that phenomena in the context of the Kermack-McKendrick model was recently discovered, showing that the overshoot is derived from a trade-off from the basic reproduction number in driving both the final epidemic size and how quickly the disease burns through the population [nguyen_fundamental_2023]. Here we lay the foundation to calculate the upper bound for overshoot when considering incidence terms beyond the simple bilinear case. We will first derive the behavior of the overshoot for more general incidence rates within the context of ODE models, where the results and understanding can be analytical and precise. As nonlinear incidence can also arise through the connectivity structure of networks, we will then numerical explore the effect of network structure and changing network topologies on overshoot.
@@ Results
$S_{t^{*}}$ $=f^{-1}(\frac{I_{t^{*}}}{g(I_{t^{*}})}\frac{1}{R_{0}})$
@@
$Overshoot=f^{-1}(\frac{1}{R_{0}})-S_{\infty}$ (10)
@@
We now turn to what restrictions there are on the form of $f(S)$. We start first with the two boundary conditions for $S$. First, we must enforce that:
@@
Otherwise, since $I$ does not have such a restriction, violating this condition leaves open the possibility of having $\frac{dS}{dt}$ (4) be non-zero when the number of susceptibles is zero, which is not realistic.
For the second boundary condition, we must have that:
$f(S=1)>\frac{\gamma}{\beta}=\frac{1}{R_{0}}$ (13)
This is obtained by inspecting $\frac{dI}{dt}$ (5) and ensuring that the incidence term $\beta f(S)g(I)$ is larger than the recovery term $\gamma I$ at $t=0$ by having a sufficiently high $\beta$. Otherwise, the epidemic never starts. Assuming $I_{0}$ is sufficiently small, then $S$ is approximately $1$ at $t=0$.
Beyond the boundary conditions, an obvious requirement is that $f(S)$ should be a continuous function. In order to be able to calculate the maximal overshoot analytically, the function should be integratable with respect to $S$ and should also have a closed-form inverse $f^{-1}$. As we will demonstrate, non-monotonic functions for $f(S)$ are possible.
@@
Conversely, there are many examples of functions that would not work. An example that satisfies the boundary conditions but that does not have a closed form inverse is $f(S)=\text{log}(S+1)$. While similar to examples listed, the following violate the boundary conditions: $exp(S),\text{log}(S)$, $\text{cos}(S)$. Examples that violate conditions of continuity include step functions of $S$ or $f(S)$ with cusps.
@@
Example: $f(S)=exp(S)-1$
Let us consider an incidence rate that takes on an exponential form. This produces an incidence term that grows slightly faster than the original bilinear incidence term, and so might be relevant in situations where there are network effects. We start at Step A. by solving for the rate of change of $I$ as a function of $S$ by taking the ratio of $\frac{dI}{dt}$ and $\frac{dS}{dt}$.
@@
from which it follows on integration using the substitution $u=e^{S}$ and partial fractions ($\frac{1}{u-1}$ and $\frac{1}{u}$) that $I+S+\frac{S-ln|e^{S}-1|}{R_{0}}$ is constant along all trajectories.
@@
$I_{0}+S_{0}+\frac{S_{0}-ln|e^{S_{0}}-1|}{R_{0}}$ $=I_{\infty}+S_{\infty}+\frac{S_{\infty}-ln|e^{S_{\infty}}-1|}{R_{0}}$
@@
$\frac{1}{R_{0}}=\frac{I_{\infty}+S_{\infty}-I_{0}-S_{0}}{S_{0}-S_{\infty}+ln(\frac{|e^{S_{\infty}}-1|}{||e^{S_{0}}-1|})}$ (15)
@@
$\frac{1}{R_{0}}$ $=\frac{S_{\infty}-1}{1-S_{\infty}+ln(\frac{|e^{S_{\infty}}-1|}{e-1})}$ (16)
@@
$Overshoot$ $=ln(\frac{ln(\frac{|e^{S_{\infty}}-1|}{e-1})}{1-S_{\infty}+ln(\frac{|e^{S_{\infty}}-1|}{e-1})})-S_{\infty}$ (18)
@@
00 0 $=(\frac{ln(\frac{|e^{S^{*}_{\infty}}-1|}{e-1})}{1-S^{*}_{\infty}+ln(\frac{|e^{S^{*}_{\infty}}-1|}{e-1})})^{-1}(\frac{(1-S^{*}_{\infty}+ln(\frac{|e^{S^{*}_{\infty}}-1|}{e-1}))\cdot((\frac{|e^{S^{*}_{\infty}}-1|}{e-1})^{-1}\frac{e^{S^{*}_{\infty}}}{e-1})-(ln(\frac{|e^{S^{*}_{\infty}}-1|}{e-1}))\cdot(-1+(\frac{|e^{S^{*}_{\infty}}-1|}{e-1})^{-1}\frac{e^{S^{*}_{\infty}}}{e-1})}{(1-S^{*}_{\infty}+ln(\frac{|e^{S^{*}_{\infty}}-1|}{e-1}))^{2}})-1$
@@
$ln(\frac{e^{S^{*}_{\infty}}-1}{e-1})(ln(\frac{e^{S^{*}_{\infty}}-1}{e-1})-S^{*}_{\infty})$ $=(1-S^{*}_{\infty})(\frac{e^{S^{*}_{\infty}}}{e^{S^{*}_{\infty}}-1})$
@@
Figure 1: The overshoot as a function of $R_{0}$ for an SIR model with nonlinear incidence term of $\beta(e^{S}-1)I$. The horizontal line for $Overshoot^{*}$ in dark blue and the vertical line in red given by $R_{0}^{*}=1.7$ are the theoretical predictions given by the calculations in the text. The curve is obtained from numerical simulations.
@@
In the previous sections, we focused on nonlinear incidence that was generated from the introduction of a nonlinearity in an ordinary differential equation model. Importantly, this fundamentally assumes homogeneity in the transmission, in that all infected individuals are identical in their ability to spread the disease further. In contrast, network models allow for heterogeneous spreading which depends on the local connectivity of each infected individual. This provides an entirely different mechanism through which nonlinear incidence can be generated compared to the ordinary differential equation models. As many real-world complexities and details can be more easily captured and explored in a network model, the consequences of those heterogeneities on the overshoot becomes more transparent in network simulations. However, a trade-off for the increased realism is that it becomes more difficult to perform analytical calculations for a general network model, so here we conduct a numerical exploration of the behavior of the overshoot in network models across a range of network structure.
@@
The space of all possible network configurations is immense, so we must restrict ourselves to analyzing a particular subset of possible networks. Here we explore what happens to the overshoot when the contact structure of the population is given by a network graph that is roughly one giant component. While it is possible to construct pathological graphs that produce very complex dynamics, we consider more classical graphs here. Using a parameterization of heterogeneity ($\sigma$) given by Ozbay et al. [ozbay_parameterizing_2023] and the configuration model of Newman [newman_networks_2018] to randomly construct networks (see Methods for details), we simulated epidemics on a spectrum of networks with structure ranging from the homogeneous limit (well-mixed, complete graph) to a heterogeneous limit (heavy-tailed degree distributions). As the dynamics of the epidemic on a network are stochastic and occur in discrete-time here, they are not parameterized by $R_{0}$ as in the ODE models. Instead, we use an analogous parameter that we will denote as a basic reproduction number for networks ($R_{0,network}$) and is defined as follows:
$R_{0,network}\equiv\frac{\tau\cdot\text{Mean Network Degree}}{\rho}$ (22)
@@
The transmission probability ($\tau$) and recovery probability ($\rho$) are directly analogous to their counterparts ($\beta$ and $\gamma$ respectively) in $R_{0}$. The inclusion of the mean network degree as a scalar makes intuitive sense as that represents the average number of potential neighbors an infected node could potentially infect at the beginning of the epidemic, which is analogous to the interpretation of $R_{0}$ as the number of secondary infections. We observed what happens to the overshoot on these different graphs as we changed $R_{0,network}$.
@@
On a network model, where contact structure is made explicit, the homogeneous limit is a complete graph, which recapitulates the well-mixed assumption of the Kermack-McKendrick model. It is not surprising then that the overshoot in the homogeneous graphs ($\sigma\approx 0)$ peaks also around 0.3 (Figure 2, $purple$ and $blue$), which coincides with the analytical upper bound previously found in the ordinary differential equation Kermack-McKendrick model [nguyen_fundamental_2023].
@@
We also see that increasing contact heterogeneity (i.e. $\sigma\rightarrow 1$) qualitatively suppresses the overshoot peak both in terms of the overshoot value and the corresponding $R_{0,network}$. Furthermore, increased heterogeneity also flattens out the overshoot curve as a function of $R_{0,network}$. We see that for more heterogeneous graphs, the overshoot is larger at very low $R_{0,network}$. For some intermediate values of contact heterogeneity (Figure 2, $yellow$ and $orange$), the overshoot shows larger overshoot than the homogeneous case for large $R_{0,network}$. This would suggest a larger public health hazard for a larger range of communicable diseases for networks with structure in this regime. The intuition for the larger overshoot at high $R_{0,network}$ in more heterogeneous networks is that both the peak occurs earlier and that a significant number of cases can occur in the periphery of a network (Figure S3-S5) Since the network core by is close to any randomly chosen node in a heterogeneous network, the core becomes more accessible the more heterogeneous the graph is which drives the peak of infections to occur sooner. In the second phase of the epidemic as the epidemic expands out to the periphery the epidemic burns more slowly as individuals in the periphery have fewer neighbors.
@@
Figure 2: The overshoot for SIR epidemic simulations on networks with varying levels of heterogeneity ($\sigma$) as a function of $R_{0,network}$. Each color shows simulations for networks of different contact heterogeneity as parameterized by $\sigma$. The solid lines represent the mean value of 100 simulation runs for a given $R_{0,network}$ and $\sigma$. The shaded areas indicate the 25th and 75th quartiles for those 100 simulations. The other simulation parameters are number of nodes (N) = 200, mean degree of network = 7, and recovery probability ($\rho$) = 0.2.

### v5

@@ Abstract
We expand the calculation of the upper bound on epidemic overshoot in SIR models to account for nonlinear incidence. We lay out the general procedure and restrictions to perform the calculation analytically for nonlinear functions in the number of susceptibles. We demonstrate the procedure by working through several examples and also numerically study what happens to the upper bound on overshoot when nonlinear incidence manifests in the form of epidemic dynamics over a contact network. We find that both steeper incidence terms and larger contact heterogeneity can increase the range of communicable diseases at which the overshoot remains a relatively large public health hazard.
@@ Introduction
Compartmental models have been an invaluable tool for analyzing the dynamics of epidemics for the last century. In particular, the susceptible-infected-recovered (SIR) model has long been a workhorse compartmental model for describing transient epidemics due to its relative simplicity and has received a lot of attention in both the academic literature and the public health arena [anderson_infectious_1992, diekmann_mathematical_2013]. The mechanism of transmission of a communicable disease underlying this model and a variety of other contagion models is contact between healthy and infectious individuals, which results in conversion of the healthy individuals into an infected state. Choosing how to precisely define the transmission interaction between healthy and infectious individuals leads to a variety of models depending on the assumptions made. For instance, in a spatially-explicit model, one might represent the contact between individual members of the population as a network. Alternatively, if one is willing to assume all individuals within a compartment are identical, one can formulate a model using ordinary differential equations (ODEs) by assuming an incidence term for how the susceptible and infected compartments mix and generate new infected individuals.
@@
Overshoot is a concept from mathematics and control theory that quantifies the amount of excess from when a function exceeds its target value [doyle_feedback_2013, sethi_what_2019]. This concept has been applied to a variety of contexts ranging from ecological and environmental problems that pertain to overconsumption and sustainability [catton_overshoot_1982, wackernagel_tracking_2002, bradshaw_underestimating_2021, fanning_social_2022] to biotechnology that controls blood flow [rosengarten_overshoot_2002]. In the epidemiological context, the overshoot quantifies the number of individuals that become infected after the prevalence peak of infections occurs (Figure 1a). In simple epidemics, as the peak of the epidemic coincides with the threshold at which transmission is sufficiently reduced so that the epidemic is no longer growing, the overshoot reflects the excess in cases beyond this minimal threshold of protection.
Figure 1: Illustration of overshoot in the context of epidemics and spatial heterogeneity in contact networks. a) In the context of an epidemic outbreak, the overshoot is given by the depletion of susceptibles from the peak of the infections until the end of the outbreak. b) Nonlinear incidence can manifest in the form of a spatial proximity network.
@@
While the overshoot within the SIR model has received some attention [handel_what_2006, ellison_implications_2020, rachel_analytical_2020, ketcheson_optimal_2021], a detailed understanding of its full mathematical behavior within compartmental models remains incomplete. An observation that the overshoot is largest for intermediate basic reproduction numbers was first numerically observed by Zarnitsyna et al. [zarnitsyna_intermediate_2018]. An explanation for that phenomena in the context of the Kermack-McKendrick model was recently discovered, showing that the overshoot is derived from a trade-off from the basic reproduction number in driving both the final epidemic size and how quickly the disease burns through the population [nguyen_fundamental_2023]. Here we lay the foundation to calculate the upper bound for overshoot when considering incidence terms beyond the simple bilinear case. We will first derive the behavior of the overshoot for more general incidence rates within the context of ODE models, where the results and understanding can be analytical and precise. As nonlinear incidence can also arise through the connectivity structure of networks (Figure 1b), we will then numerical explore the effect of network structure and changing network topologies on overshoot.
@@ Results
$S_{t^{*}}$ $=f^{-1}\left(\frac{I_{t^{*}}}{g(I_{t^{*}})}\frac{1}{R_{0}}\right)$
@@
$Overshoot=f^{-1}\left(\frac{1}{R_{0}}\right)-S_{\infty}$ (10)
@@
We now turn to what restrictions there are on the form of $f(S)$. We start first with two conditions. First, we must enforce that when there are no susceptibles ($S=0$), then the incidence rate must go to zero ($\beta f(S)I=0$). Otherwise, since $I$ does not have such a restriction, violating this condition would leave open the unrealistic possibility that the model can generate infected people when there are no susceptibles available. To ensure this condition is met, we need the function on S to output zero if the input is zero.
@@
For the second condition, for an outbreak to occur in an SIR model, we must have a minimum value for the basic reproduction number, $R_{0}$. Another way to view this condition is by inspecting (2) and recognizing that that the incidence term ($\beta f(S)I$) needs to be greater than the recovery term ($\gamma I$). Otherwise the epidemic cannot grow in size. Comparing the two terms leaves an inequality ($\beta f(S)I>\gamma I$) which can be rewritten as:
$R_{0}>\frac{1}{f(S)}$ (13)
Beyond these conditions, an obvious requirement is that $f(S)$ should be a continuous function. In order to be able to calculate the maximal overshoot analytically, the function should be integratable with respect to $S$ and should also have a closed-form inverse $f^{-1}$. As we will demonstrate, non-monotonic functions for $f(S)$ are possible.
@@
Conversely, there are many examples of functions that would not work. An example that satisfies the boundary conditions but that does not have a closed form inverse is $f(S)=\text{log}(S+1)$. While similar to examples listed, the following violate the boundary conditions: $\text{exp}(S),\text{log}(S)$, $\text{cos}(S)$. Examples that violate conditions of continuity include step functions of $S$ or $f(S)$ with cusps.
@@
Example: $f(S)=\text{exp}(S)-1$
Let us consider an incidence rate that takes on an exponential form. This produces an incidence term that grows slightly faster than the original bilinear incidence term, and so might be relevant in situations where there are network effects.The phenomenon of superspreading occurs when some individuals infect many more people than the typical infected person would normally do [majra_sars-cov-2_2021]. Superspreading allows for explosive outbreaks that exceed the growth rate of infections allowed by simple bilinear incidence and is enabled by inherent heterogeneity in contacts amongst the population, which makes it natural to occur in situations that are approximated by a network. Different models can be formulated to account for superspreading [fujie_effects_2007, nielsen_covid-19_2021]. Here we have chosen the exponential as a qualitative simple example that more closely resembles the growth rate on a heterogeneous network than standard bilinear incidence.
@@
from which it follows on integration using the substitution $u=e^{S}$ and partial fractions $\left(\frac{1}{u-1}\right)$ and $\left(\frac{1}{u}\right)$ that $I+S+\frac{S-\text{ln}|e^{S}-1|}{R_{0}}$ is constant along all trajectories.
@@
$I_{0}+S_{0}+\frac{S_{0}-\text{ln}|e^{S_{0}}-1|}{R_{0}}$ $=I_{\infty}+S_{\infty}+\frac{S_{\infty}-\text{ln}|e^{S_{\infty}}-1|}{R_{0}}$
@@
$\frac{1}{R_{0}}=\frac{I_{\infty}+S_{\infty}-I_{0}-S_{0}}{S_{0}-S_{\infty}+\text{ln}\left(\frac{|e^{S_{\infty}}-1|}{||e^{S_{0}}-1|}\right)}$ (15)
@@
$\frac{1}{R_{0}}$ $=\frac{S_{\infty}-1}{1-S_{\infty}+\text{ln}\left(\frac{|e^{S_{\infty}}-1|}{e-1}\right)}$ (16)
@@
$Overshoot$ $=\text{ln}\left(\frac{\text{ln}\left(\frac{|e^{S_{\infty}}-1|}{e-1}\right)}{1-S_{\infty}+\text{ln}\left(\frac{|e^{S_{\infty}}-1|}{e-1}\right)}\right)-S_{\infty}$ (18)
@@
00 0 $=\left(\frac{\text{ln}\left(\frac{|e^{S^{*}_{\infty}}-1|}{e-1}\right)}{1-S^{*}_{\infty}+\text{ln}\left(\frac{|e^{S^{*}_{\infty}}-1|}{e-1}\right)}\right)^{-1}\left(\frac{\left(1-S^{*}_{\infty}+\text{ln}\left(\frac{|e^{S^{*}_{\infty}}-1|}{e-1}\right)\right)\cdot\left(\left(\frac{|e^{S^{*}_{\infty}}-1|}{e-1}\right)^{-1}\frac{e^{S^{*}_{\infty}}}{e-1}\right)-\text{ln}\left(\frac{|e^{S^{*}_{\infty}}-1|}{e-1}\right))\cdot\left(-1+\left(\frac{|e^{S^{*}_{\infty}}-1|}{e-1}\right)^{-1}\frac{e^{S^{*}_{\infty}}}{e-1}\right)}{\left(1-S^{*}_{\infty}+\text{ln}\left(\frac{|e^{S^{*}_{\infty}}-1|}{e-1}\right)\right)^{2}}\right)-1$
@@
$\text{ln}\left(\frac{e^{S^{*}_{\infty}}-1}{e-1}\right)\left(\text{ln}\left(\frac{e^{S^{*}_{\infty}}-1}{e-1}\right)-S^{*}_{\infty}\right)$ $=(1-S^{*}_{\infty})\left(\frac{e^{S^{*}_{\infty}}}{e^{S^{*}_{\infty}}-1}\right)$
@@
Figure 2: The overshoot in a model with exponential incidence. The overshoot as a function of $R_{0}$ for an SIR model with nonlinear incidence term of $\beta(e^{S}-1)I$. The horizontal line for $Overshoot^{*}$ in dark blue and the vertical line in red given by $R_{0}^{*}=1.7$ are the theoretical predictions given by the calculations in the text. The curve is obtained from numerical simulations.
@@
In the previous sections, we focused on nonlinear incidence that was generated from the introduction of a nonlinearity in an ordinary differential equation model. Importantly, this fundamentally assumes homogeneity in the transmission, in that all infected individuals are identical in their ability to spread the disease further. In contrast, network models allow for heterogeneous spreading which depends on the local connectivity of each infected individual. This provides an entirely different mechanism through which nonlinear incidence can be generated compared to the ordinary differential equation models. As many real-world complexities and details can be more easily captured and explored in a network model, the consequences of those heterogeneities on the overshoot becomes more transparent through the targeted and fully-fleshed out explorations that can be done through numerical experimentation in network simulations [newman_spread_2002, wang_epidemic_2003, keeling_networks_2005, pastor-satorras_epidemic_2015, kiss_mathematics_2017]. However, a trade-off for the increased realism is that it becomes more difficult to perform analytical calculations for a general network model, so here we conduct a numerical exploration of the behavior of the overshoot in network models across a range of network structure.
@@
The space of all possible network configurations is immense, so we must restrict ourselves to analyzing a particular subset of possible networks. Here we explore what happens to the overshoot when the contact structure of the population is given by a network graph that is roughly one giant component. While it is possible to construct pathological graphs that produce very complex dynamics, we consider more classical graphs here. Using a parameterization of heterogeneity ($\sigma$) given by Ozbay et al. [ozbay_parameterizing_2023] and the configuration model of Newman [newman_networks_2018] to randomly construct networks (see Methods for details), we simulated epidemics on a spectrum of networks with structure ranging from the homogeneous limit (well-mixed, complete graph) to a heterogeneous limit (heavy-tailed degree distributions). The spectrum of distributions shapes across the space of $\sigma$ used to generate the degree distributions can be seen in the supplement. As the dynamics of the epidemic on a network are stochastic and occur in discrete-time here, they are not parameterized by $R_{0}$ as in the ODE models. Instead, we use an analogous parameter that we will denote as a basic reproduction number for networks ($R_{0,network}$) and is defined as follows:
$R_{0,network}\equiv\frac{\tau\cdot\langle k\rangle}{\rho}$ (21)
@@
The transmission probability ($\tau$) and recovery probability ($\rho$) are directly analogous to their counterparts ($\beta$ and $\gamma$ respectively) in $R_{0}$ in that that correspond to transmission and recovery parameters, albeit for a stochastic model. The inclusion of the mean network degree ($\langle k\rangle$) as a scalar makes intuitive sense as that represents the average number of potential neighbors an infected node could potentially infect at the beginning of the epidemic, which is analogous to the interpretation of $R_{0}$ as the number of secondary infections. We observed what happens to the overshoot on these different graphs as we changed $R_{0,network}$.
@@
On a network model, where contact structure is made explicit, the homogeneous limit is a complete graph, which recapitulates the well-mixed assumption of the Kermack-McKendrick model. It is not surprising then that the overshoot in the homogeneous graphs ($\sigma\approx 0)$ peaks also around 0.3 (Figure 3, $purple$ and $blue$), which coincides with the analytical upper bound previously found in the ordinary differential equation Kermack-McKendrick model [nguyen_fundamental_2023]. However, while the homogeneous graphs shown here are highly regular (i.e. symmetric in the average number of contacts each node has), the network’s connectivity is not close to complete as the mean degree is significantly less than the network size. Thus, the average overshoot is lower than would be the case for a complete graph.
@@
We also see that increasing contact heterogeneity (i.e. $\sigma\rightarrow 1$) qualitatively suppresses the overshoot peak both in terms of the overshoot value and the corresponding $R_{0,network}$. Furthermore, increased heterogeneity also flattens out the overshoot curve as a function of $R_{0,network}$. We see that for more heterogeneous graphs, the overshoot is larger at very low $R_{0,network}$. For some intermediate values of contact heterogeneity (Figure 3, $yellow$ and $orange$), the overshoot shows larger overshoot than the homogeneous case for large $R_{0,network}$. This would suggest a larger public health hazard for a larger range of communicable diseases for networks with structure in this regime. The intuition for the larger overshoot at high $R_{0,network}$ in more heterogeneous networks is that both the peak occurs earlier and that a significant number of cases can occur in the periphery of a network (Supplemental Materials). The probability of being connected to a high-degree node increases as the heterogeneity of the network increases, which drives the peak of infections to occur sooner. In the second phase of the epidemic as the epidemic expands out to the periphery the epidemic burns more slowly as individuals in the periphery have fewer neighbors.
@@
Figure 3: Comparing the overshoot in the context of networks with different amounts of contact heterogeneity. The overshoot for SIR epidemic simulations on networks with varying levels of heterogeneity ($\sigma$) as a function of $R_{0,network}$. Each color shows simulations for networks of different contact heterogeneity as parameterized by $\sigma$. The solid lines represent the mean value of 100 simulation runs for a given $R_{0,network}$ and $\sigma$. The shaded areas indicate the 25th and 75th quartiles for those 100 simulations. The other simulation parameters are number of nodes (N) = 200, $\langle k\rangle$ = 7, and recovery probability ($\rho$) = 0.2.